OpenStudy (anonymous):

From a group of 7 men and 6 women, five persons are to be selected to form a committee so that at least 3 men are there on the committee. In how many ways can it be done?

Parth (parthkohli):

I think it is\[\left[\binom{7}{3}\times \binom{6}{2}\right] + \left[\binom{7}{4} \times \binom{6}{1}\right] + \left[\binom{7}{5}\right]\]

Parth (parthkohli):

Since the possible ways are: 3 men and 2 women 4 men and 1 woman 5 men

terenzreignz (terenzreignz):

I prefer \[\Large \left(\begin{matrix}7\\ 3\end{matrix}\right)\times \left(\begin{matrix}10\\ 2\end{matrix}\right)\] 3 men and 2 more of heaven-knows-what-gender

OpenStudy (anonymous):

Whom do i prefer? :P

terenzreignz (terenzreignz):

I don't know... :D

OpenStudy (anonymous):

ohk.. @terenzreignz

terenzreignz (terenzreignz):

Well... @ParthKohli ?

Parth (parthkohli):

Hmm, our answers are different. But I'm not able to get who has their answer incorrect lol

terenzreignz (terenzreignz):

ehh.. stick to @ParthKohli I probably oversimplified to ... oblivion :D

terenzreignz (terenzreignz):

Today is not my day o.O

Parth (parthkohli):

It's not mine either.

OpenStudy (agent0smith):

@terenzreignz your answer is def not correct :P

terenzreignz (terenzreignz):

That settles it.

OpenStudy (agent0smith):

Hmm, @terenzreignz now the more i think about it, the more I wonder why your answer doesn't work... it seems reasonable, but gives over double the possible combinations as parth's method. Yours guarantees having at least 3 men... but I'm guessing it counts a few situations twice. ie you pick Joe, Barry and Larry from the first 7, then Harry and Sam from the second group of 10. Then you could also pick Joe, sam and harry from the first 7, then also pick Barry and Larry from the second group of 10, etc.

terenzreignz (terenzreignz):

I've already put it out of my mind, unfortunately :D But Parth's method seems so fail-safe...so

OpenStudy (agent0smith):

His method avoids possibly over-counting the same group twice, or thrice, or... four... ice? Frice?

OpenStudy (anikhalder):

Possible groupings :- 1. 3 men and 2 women 2. 4 men and 1 woman 3. 5 men These 3 results are independent of each other, hence we got to calculate each individually and add the 3 results. No of ways to choose 3 men from 7 and 2 women from 6 = 7C3 x 6C2=525 No of ways to choose 2 men from 7 and 1 women from 6 = 7C2 x 6C1=126 No of ways to choose 5 men from 7 and 0 women from 6 = 7C5 x 6C0=21 Total no. of ways= 525+126+21= 672 (This is a problem on Combinations) I hope you are familiar with Permutations and Combinations?

OpenStudy (agent0smith):

No of ways to choose 2 men from 7 and 1 women from 6 = 7C2 x 6C1=126 @anikhalder that should be choosing 4 men.

OpenStudy (anikhalder):

oh yea..i'm sorry..it would be 4..thanks @agent0smith and 1 woman instead of 1 women..:) 7C4 x 6C1= 210 So total is= 525+210+21=756